OpenStudy (joachim):

Suppose that f is an even function and g is an odd function. Are the following functions odd, even or neither?

OpenStudy (joachim):

h(x) = f(x) + g(x) k(x) = f(x) * g(x) l(x) = f(g(x))

OpenStudy (joachim):

I would have said in the following order: neither, odd, even... Is that correct?

OpenStudy (anonymous):

neither for first one

OpenStudy (anonymous):

odd for second one, so far so good

OpenStudy (anonymous):

even for last one three for three!